Circular No. 6246

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN 3 The Bureau has received numerous reports from observers worldwide of

independent discoveries of this comet, which has continued its remarkable outburst in brightness (cf. IAUC 6234), as shown by the following total magnitude estimates (B = binoculars): Sept. 30.12 UT, 7.4 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); Oct. 2.02, 6.6 (G. W. Kronk, Troy, IL, 20×80 B); 7.42, 6.5 (G. Garradd, Loomberah, N.S.W., 10×50 B); 9.91, 5.8 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 B); 11.39, 5.8 (T. Lovejoy, Jimboomba, Queensland, Australia, 15×80 B; 1^{o} dust tail); 13.07, 5.4 (A. Hale, Cloudcroft, NM, 10×50 B); 15.39, 5.7 (Lovejoy; 1^{o}.5 tail); 17.01, 6.0 (Kronk); 19.06, 5.9 (R. Keen, Mt. Thorodin, CO, 7×50 B).

Ephemeris from the orbital elements on MPC 25030:

1995TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 10 15^{h}57.^{m}96 −26^{o}13.^{′}4 1.320 0.966 46.^{o}7 48.^{o}8 5.5
15 16 25.32 −28 11.8 1.314 0.988 48.2 48.9 5.5
20 16 53.72 −29 49.3 1.314 1.014 49.8 48.6 5.7
25 17 22.85 −31 03.2 1.322 1.044 51.3 48.0 5.8
30 17 52.29 −31 51.7 1.336 1.078 52.7 47.1 6.0
Nov. 4 18 21.58 −32 14.3 1.357 1.114 54.0 46.0 6.1
9 18 50.27 −32 12.1 1.385 1.154 55.1 44.8 6.3
14 19 17.96 −31 47.2 1.421 1.195 56.1 43.4 6.5
19 19 44.36 −31 02.5 1.463 1.238 56.8 41.9 6.8
24 20 09.29 −30 01.6 1.510 1.282 57.3 40.4 7.0
29 20 32.66 −28 47.8 1.564 1.327 57.5 38.8 7.2

SUPERNOVA 1995ag IN UGC 11861 A. V. Filippenko and A. J. Barth, University of California at Berkeley,

write: “CCD spectrograms (range 310–1000 nm, resolution 0.7–1.4 nm) obtained on Oct. 17 UT with the 3-m Shane reflector at Lick Observatory show that SN 1995ag is actually of type II, not type I (cf. IAUC 6244). There are prominent, broad hydrogen Balmer lines having P-Cyg profiles.”

P. Sicoli reports the following precise position of SN 1995ag from an observation made at Sormano, Italy, on Oct. 16.828 UT by F. Manca and V. Giuliani (0.5-m reflector + CCD): α = 21^{h}56^{m}19^{s}.4, δ = +73^{o}15′13′′.6 (equinox 2000.0).

1995 October 19 Daniel W. E. Green